1.5 molar,




50 mL nitric acid solution



Volume of solute, mass solute, volume of water

1
Expert's answer
2022-04-25T20:47:04-0400

C = n(solute) / V(solution)

n(solute) = V(solution) × C = 0.05 L × 1.5 mol/L = 0.075 mol

m(solute) = n(solute) × Mr(solute) = 0.075 mol × 63.01 g/mol (HNO3) = 4.73 g

V(HNO3) = m(HNO3) / "\rho"(HNO3) = 4.73 g / 1.51 g/mL = 3.13 mL

V(water) = V(solution) - V(HNO3) = 50 mL - 3.13 mL = 46.87 mL